# Magnon topology driven by altermagnetism

### v1

@@ Abstract
Altermagnets present a new class of fully compensated collinear magnetic order, where the two sublattices are not related merely by time-reversal combined with lattice translation or inversion, but require an additional lattice rotation. This distinctive symmetry leads to a characteristic splitting of the magnon bands; however the splitting is only partial – residual degeneracies persist along certain lines in the Brillouin zone as a consequence of the underlying altermagnetic rotation. We consider a two-dimensional $d$-wave altermagnetic spin model on the checkerboard lattice and introduce additional interactions such as an external magnetic field and Dzyaloshinskii-Moriya interactions, that lift these degeneracies. The resulting magnon bands become fully gapped and acquire non-trivial topology, characterized by nonzero Chern numbers. We demonstrate the crucial role of altermagnetism for the generation of the Berry curvature. As a direct consequence of the topological magnons, we find finite thermal Hall conductivity $\kappa_{xy}$, which exhibits a characteristic low-temperature scaling, $\kappa_{xy}\propto T^{4}$. Moreover, $\kappa_{xy}$ changes sign under reversal of the magnetic field, exhibiting a sharp jump across zero field at low temperatures. We also demonstrate topologically protected chiral edge modes in a finite strip geometry.
@@ I Introduction
In this article, we investigate topological magnons in a minimal two-dimensional altermagnetic spin model – the checkerboard lattice spin model [57, 54, 55] in the presence of an out-of-plane uniform Zeeman field and DMI. The altermagnetic checkerboard lattice spin model consists of two exchange couplings – antiferromagnetic Heisenberg coupling between nearest-neighbors and ferromagnetic or antiferromagnetic Heisenberg coupling between next-nearest-neighbors along the diagonals of the checkerboard plaquettes. The latter is crucial for realizing altermagnetism in this model [54, 55]; in its absence, this model reduces to the nearest-neighbor square-lattice AFM. Accordingly, we refer to the diagonal coupling henceforth as the AM coupling. This checkerboard lattice model stabilizes a fully compensated collinear altermagntic ground state, with its two magnon energy-branches displaying $d$-wave-like splitting, remaining degenerate only along certain lines in the Brillouin zone [54, 55]. We find that introducing a uniform out-of-plane Zeeman field in this model stabilizes a spin-flop phase: spins cant out of the lattice-plane towards the field, and moreover, this canted state remains the ground state even upon adding the specific DMI that we consider. Using linear spin-wave theory, we analytically show that when all three ingredients – AM coupling, Zeeman field, and DMI – are present, two magnon branches become fully gapped. Interestingly, the energy gap between the bands at the mid-points of the Brillouin zone edges is determined solely by the AM coupling. To reveal the topological character of the magnon bands, we calculate the Berry connection, Berry curvature, and find the Chern number to be non-zero, namely ±$\pm 1$, and opposite for the upper and lower branches. Furthermore, reversing the sign of any of the three couplings – Zeeman field, AM coupling, or DMI, reverses the sign of the Berry curvature and hence the Chern number. In the limit of vanishing AM coupling, the Berry curvature becomes concentrated near the midpoints of the Brillouin zone edges, forming elliptical regions whose principal axes are set by the DMI and Zeeman field, respectively. As a consequence of the topological magnons, we compute thermal Hall conductivity, $\kappa_{xy}$, under a temperature gradient. At low temperatures, $\kappa_{xy}$ is found to scale with temperature $T$ as $\kappa_{xy}\propto T^{4}$. Moreover, $\kappa_{xy}$ changes sign upon flipping the Zeeman field, displaying a sharp jump across zero field at low temperatures. Finally, as another non-trivial manifestation of magnon band topology, we also demonstrate the existence of chiral edge states on a finite strip geometry.
@@ II Model
[…]
@@ III Classical ground states
$\bm{S}_{\bm{r},1}$ $=S\left(\sin\theta_{1}\cos\phi_{1}\,\bm{\hat{x}}+\sin\theta_{1}\sin\phi_{1}\,\bm{\hat{y}}+\cos\theta_{1}\,\bm{\hat{z}}\right),$
$\bm{S}_{\bm{r}^{\prime},2}$ $=S\left(\sin\theta_{2}\cos\phi_{2}\,\bm{\hat{x}}+\sin\theta_{2}\sin\phi_{2}\,\bm{\hat{y}}+\cos\theta_{2}\,\bm{\hat{z}}\right),$ (2)
@@
$\bm{S}_{\bm{r},1}\equiv\bm{S}_{1}$ $=$ $S(\cos\xi\,\bm{\hat{x}}+\sin\xi\,\bm{\hat{z}}),$
$\bm{S}_{\bm{r}^{\prime},2}\equiv\bm{S}_{2}$ $=$ $S(-\cos\xi\,\bm{\hat{x}}+\sin\xi\,\bm{\hat{z}}).$ (3)
@@ IV Spin-wave theory and magnon bands
${\cal H}_{2}\!=\!\!\sum_{\alpha\beta,\bm{q}}\left[{A}^{\alpha\beta}_{\bm{q}}{a}^{\dagger}_{\bm{q},\alpha}{a}^{\phantom{\dagger}}_{\bm{q},\beta}+\frac{1}{2}\left({B}^{\alpha\beta}_{\bm{q}}{a}^{\dagger}_{\bm{q},\alpha}{a}^{\dagger}_{-\bm{q},\beta}+\bar{B}^{\alpha\beta}_{\bm{q}}{a}^{\phantom{\dagger}}_{-\bm{q},\beta}{a}^{\phantom{\dagger}}_{\bm{q},\alpha}\right)\right],$ (4)
@@
Figure 2: Linear spin-wave spectrum with the Zeeman field $H=H_{s}/8$ and DMI constant $D=0.5J$ for a path in the first Brillouin zone ${\rm\Gamma}\,(0,0)\rightarrow{\rm X}\,(\pi/a_{0},0)\rightarrow{\rm M}\,(\pi/a_{0},\pi/a_{0})\rightarrow{\rm\Gamma}\,(0,0)\rightarrow{\rm Y}\,(0,\pi/a_{0})$. (a) Spectrum with the AM coupling $J_{1}=0$ shows degeneracy at X and Y, marked by the red circle and (b) with $J_{1}=-0.04J$, a gap opens up at X and Y.
@@
$\hbar\omega^{\pm}_{\bm{q}}=\frac{1}{\sqrt{2}}\sqrt{\left(A^{11}_{\bm{q}}\right)^{2}+\left(A^{22}_{\bm{q}}\right)^{2}-2\left(B^{12}_{\bm{q}}\right)^{2}+2\left|A^{12}_{\bm{q}}\right|^{2}\pm\mathcal{E}^{2}_{\bm{q}}}\,,$ (6)
@@
The upper magnon branch is gapped at the zone center with energy $E^{+}_{\Gamma}=|g\mu_{\textsc{b}}\mu_{0}H|$. The energy gap between the two branches at $\bm{q}={\rm{M}}\equiv(\pi/a_{0},\pi/a_{0})$ is set by a combination of DMI and the Zeeman field: $\Delta_{\rm M}=g\mu_{B}\mu_{0}|DH|/J$. In contrast, at $\bm{q}={\rm{X}}\equiv(\pi/a_{0},0)$ or $\bm{q}={\rm{Y}}\equiv(0,\pi/a_{0})$, the gap is set only by the AM coupling: $\Delta_{\rm X}=4|J_{1}|S$ [see Fig. 2(b)]. Thus, the gap at X and Y is controlled by altermagnetism, while the gap at M is determined by the combined action of DMI and the Zeeman field. The same band gap at X and Y is also a consequence of the four-fold rotation symmetry of the spectrum discussed in the previous paragraph. In this work, we will focus more on the coupling parameter regime where the gap at M is already opened by the DMI and Zeeman field, and the AM coupling is tuned from zero to a finite value – highlighting the role of altermagnetism in opening up the gap between the magnon branches at X and Y. The regime of interest in this context would be $\Delta_{\rm X}\ll\Delta_{\rm M}$.
@@ V Topology of the magnon bands: Berry curvature and Chern numbers
Figure 3: Berry connections ${\cal\bm{A}}^{\pm}_{\bm{q}}$ of the upper and lower branches for the AM coupling $J_{1}=-0.04J$, DMI constant $D=0.5J$, and Zeeman field $H=H_{s}/8$.
@@
and show it across the first Brillouin zone for the two branches in Fig. 4 for different values of the AM coupling. The $C_{4}$ lattice-rotation symmetry of the Hamiltonian [Eq. (1)] is reflected in the Berry curvature as $\Omega^{\pm}_{C_{4}\bm{q}}=\Omega^{\pm}_{\bm{q}}$, as evident in Fig. 4. Applying this symmetry twice successively results in the even-under-inversion property of the Berry curvature, i.e., $\Omega^{\pm}_{-\bm{q}}=\Omega^{\pm}_{\bm{q}}$. Finally, the topology of the magnon bands in two-dimensional systems is characterized by the (first) Chern number defined in terms of the Berry curvature as
[…]
where the integration is performed over the first Brillouin zone (FBZ). Evaluating this expression for the two gapped bands yields $|C^{\pm}|=1$ with $C^{+}=-C^{-}$, thereby clearly demonstrating that the magnon bands in our model are topologically non-trivial.
@@
Figure 4: Evolution of the Berry curvature of the upper (top row) and lower (bottom row) magnon branches as the absolute value of the AM coupling $J_{1}$ decreases from left to right. The other parameters are fixed at the DMI constant $D=0.5J$ and the Zeeman field $H=H_{s}/8$.
We find the following interesting properties of the Berry curvature and Chern number for each magnon branch: (i) Under reversal of the Zeeman field, $H\to-H$, the Berry curvature changes as $\Omega^{\pm}_{\bm{q}}\to-\Omega^{\pm}_{\bm{q}}$, resulting in $C^{\pm}\to-C^{\pm}$; (ii) Similarly, reversing the DMI, $D\to-D$, leads to $\Omega^{\pm}_{\bm{q}}\to-\Omega^{\pm}_{\bm{q}}$ and $C^{\pm}\to-C^{\pm}$; (iii) When the AM coupling is reversed, $J_{1}\to-J_{1}$, the Berry curvature also changes sign at each wave vector, but unlike the previous two cases, its absolute value does not remain unchanged. Nevertheless, the change happens in such a way that we still find $C^{\pm}\to-C^{\pm}$.
@@ VI Thermal Hall conductivity
As a physical consequence of the nontrivial topology of the magnon bands, we consider the magnon thermal Hall effect, i.e., the effect of heat transfer by magnons in the direction perpendicular to the applied temperature gradient [19, 20, 21]. This phenomenon has been studied extensively in both ferromagnetic [19, 21, 64, 22, 23, 24, 65, 31, 32, 33, 35, 39, 40, 41, 42] and antiferromagnetic [61, 43, 44, 45, 46, 47, 48, 50, 51, 66, 49] systems, while its realization in an altermagnetic system has only recently begun to attract attention [67]. To compute the thermal Hall conductivity, we use the formula [19, 20, 64]
@@
The Zeeman field dependence of $\kappa_{xy}$ is shown in Fig. 6. As the field is reversed, $\kappa_{xy}$ changes sign and, notably, exhibits a sharp jump across zero at low temperatures. It qualitatively reproduces the corresponding field dependence observed experimentally for Lu2V2O7 [21] and predicted theoretically [64, 68, 31]. This non-monotonic behavior of $\kappa_{xy}$ for low fields at low temperatures follows from the field dependence of the Berry curvature $\Omega^{-}_{\bm{q}}$ of the lower magnon branch near $\Gamma$ point, see the inset in Fig. 6. Note that the magnons in the lower branch from the vicinity of $\Gamma$ point mainly contribute to $\kappa_{xy}$ at low temperatures.

### v2

@@ Abstract
Altermagnets present a class of fully compensated collinear magnetic order, where the two sublattices are not related merely by time-reversal combined with lattice translation or inversion, but require an additional lattice rotation. This distinctive symmetry leads to a characteristic splitting of the magnon bands; however the splitting is only partial – residual degeneracies persist along certain lines in the Brillouin zone as a consequence of the underlying altermagnetic rotation. We consider a two-dimensional $d$-wave altermagnetic spin model on the checkerboard lattice and introduce additional interactions such as an external magnetic field and Dzyaloshinskii-Moriya interactions, that lift these degeneracies. The resulting magnon bands become fully gapped and acquire non-trivial topology, characterized by nonzero Chern numbers. We demonstrate the crucial role of altermagnetism for the generation of the Berry curvature. As a direct consequence of the topological magnons, we find finite thermal Hall conductivity $\kappa_{xy}$, which exhibits a characteristic low-temperature scaling, $\kappa_{xy}\propto T^{4}$. Moreover, $\kappa_{xy}$ changes sign under reversal of the magnetic field, exhibiting a sharp jump across zero field at low temperatures. We also demonstrate topologically protected chiral edge modes in a finite strip geometry.
@@ I Introduction
In this article, we investigate topological magnons in a minimal two-dimensional altermagnetic spin model – the checkerboard lattice spin model [57, 54, 55] in the presence of an out-of-plane uniform Zeeman field and DMI. The altermagnetic checkerboard lattice spin model consists of two exchange couplings – antiferromagnetic Heisenberg coupling between nearest-neighbors and ferromagnetic or antiferromagnetic Heisenberg coupling between next-nearest-neighbors along the diagonals of the checkerboard plaquettes. The latter is crucial for realizing altermagnetism in this model [54, 55]; in its absence, this model reduces to the nearest-neighbor square-lattice AFM. Accordingly, we refer to the diagonal coupling henceforth as the AM coupling. This checkerboard lattice model stabilizes a fully compensated collinear altermagnetic ground state, with its two magnon energy-branches displaying $d$-wave-like splitting, remaining degenerate only along certain lines in the Brillouin zone [54, 55]. We find that introducing a uniform out-of-plane Zeeman field in this model stabilizes a spin-flop phase: spins cant out of the lattice-plane towards the field, and moreover, this canted state remains the ground state even upon adding the specific DMI that we consider. Using linear spin-wave theory, we analytically show that when all three ingredients – AM coupling, Zeeman field, and DMI – are present, two magnon branches open a gap with respect to each other. Interestingly, the energy gap between the bands at the mid-points of the Brillouin zone edges is determined solely by the AM coupling. To reveal the topological character of the magnon bands, we calculate the Berry connection, Berry curvature, and find the Chern number to be non-zero, namely $\pm 1$, and opposite for the upper and lower branches. Furthermore, reversing the sign of any of the three couplings – Zeeman field, AM coupling, or DMI, reverses the sign of the Berry curvature and hence the Chern number. In the limit of vanishing AM coupling, the Berry curvature becomes concentrated near the midpoints of the Brillouin zone edges, forming elliptical regions whose principal axes are set by the DMI and Zeeman field, respectively. As a consequence of the topological magnons, we compute thermal Hall conductivity, $\kappa_{xy}$, under a temperature gradient. At low temperatures, $\kappa_{xy}$ is found to scale with temperature $T$ as $\kappa_{xy}\propto T^{4}$. Moreover, $\kappa_{xy}$ changes sign upon flipping the Zeeman field, displaying a sharp jump across zero field at low temperatures. Finally, as another non-trivial manifestation of magnon band topology, we also demonstrate the existence of chiral edge states on a finite strip geometry.
@@ II Model
${\cal H}$ $=J\sum_{\langle\bm{r}\bm{r}^{\prime}\rangle}\bm{S}_{\bm{r},\,1}\cdot\bm{S}_{\bm{r}^{\prime},\,2}+J_{1}\sum_{\bm{r}}\bm{S}_{\bm{r},\,1}\cdot\bm{S}_{\bm{r}\,+\,a_{0}\bm{\hat{y}},\,1}$
$+J_{1}\sum_{\bm{r}^{\prime}}\bm{S}_{\bm{r}^{\prime},\,2}\cdot\bm{S}_{\bm{r}^{\prime}\,+\,a_{0}\bm{\hat{x}},\,2}-g\mu_{\textsc{b}}\mu_{0}\bm{H}\cdot\bigg[\!\sum_{\bm{r}}\bm{S}_{\bm{r},\,1}+\sum_{\bm{r}^{\prime}}\bm{S}_{\bm{r}^{\prime},\,2}\bigg]$
$+D\bm{\hat{z}}\cdot\!\!\sum_{\bm{r}}\bm{S}_{\bm{r},\,1}\times\Big[\bm{S}_{\bm{r}\,+\,\bm{\delta}_{\searrow},\,2}+\bm{S}_{\bm{r}\,-\,\bm{\delta}_{\searrow},\,2}-\bm{S}_{\bm{r}\,+\,\bm{\delta}_{\nearrow},\,2}-\bm{S}_{\bm{r}\,-\,\bm{\delta}_{\nearrow},\,2}\Big],$ (1)
@@ III Classical ground states
[…]
where $\{\bm{\hat{x}},\bm{\hat{y}},\bm{\hat{z}}\}$ constitutes a right-handed laboratory coordinate frame [see Fig. 1], and $\theta_{\alpha}$ and $\phi_{\alpha}$ with $\alpha=1,\,2$ are the polar and azimuthal angles specifying the classical spin-direction on sublattice-$\alpha$. Using these spin vectors, and assuming the spatially uniform magnetization of each of the sublatices, we compute the energy from Eq. (1) and find the following: since the AM coupling ($J_{1}$) terms involve spins only within the same sublattice, they do not depend on $\theta_{\alpha}$ or $\phi_{\alpha}$. Moreover, the total DMI for the four nearest-neighbor bonds of any sublattice sums to zero because the total Dzyaloshinskii-Moriya vector of a magnetic unit cell vanishes (see Fig. 1). Consequently, the ground state energy does not depend on DMI. Therefore, determining $\theta_{\alpha}$ and $\phi_{\alpha}$ depends only on $J$ and the Zeeman field, reducing the problem to finding the ground-state configuration of a nearest-neighbor Heisenberg AFM on the square lattice in the presence of a uniform Zeeman field [58]. The ground state undergoes a coplanar spin-flop phase where the two sublattices cant out of the lattice plane towards the field direction by an angle $\xi=\sin^{-1}\left(\frac{g\mu_{\textsc{b}}\mu_{0}H}{8JS}\right)$ with the lattice plane. This yields $\theta_{1}=\theta_{2}=\pi/2-\xi$ and $\phi_{1}-\phi_{2}=\pi$. At $H=0$, we have $\xi=0$ corresponding to a Néel state in the lattice plane. As $H$ increases, two sublattices progressively cant towards the field direction. At the saturation field, ${H}_{s}=\frac{8JS}{g\mu_{\textsc{b}}\mu_{0}}$, we have $\xi=\pi/2$, and the system reaches a fully polarized configuration along the field direction. As a general reference ground state, we may choose $\phi_{1}=0$ and $\phi_{2}=\pi$, such that the ground state is given by
[…]
@@ IV Spin-wave theory and magnon bands
$\hskip-2.84544pt{\cal H}_{2}\!=\!\!\sum_{\alpha\beta,\bm{q}}\left[{A}^{\alpha\beta}_{\bm{q}}{a}^{\dagger}_{\bm{q},\alpha}{a}^{\phantom{\dagger}}_{\bm{q},\beta}+\frac{1}{2}\left({B}^{\alpha\beta}_{\bm{q}}{a}^{\dagger}_{\bm{q},\alpha}{a}^{\dagger}_{-\bm{q},\beta}+\bar{B}^{\alpha\beta}_{\bm{q}}{a}^{\phantom{\dagger}}_{-\bm{q},\beta}{a}^{\phantom{\dagger}}_{\bm{q},\alpha}\right)\right],$ (4)
@@
Figure 2: Linear spin-wave spectrum [Eq. (6)] with the Zeeman field $H=H_{s}/8$ and DMI constant $D=0.5J$ for a path in the first Brillouin zone ${\rm\Gamma}\,(0,0)\rightarrow{\rm X}\,(\pi/a_{0},0)\rightarrow{\rm M}\,(\pi/a_{0},\pi/a_{0})\rightarrow{\rm\Gamma}\,(0,0)\rightarrow{\rm Y}\,(0,\pi/a_{0})$. (a) Spectrum with the AM coupling $J_{1}=0$ shows degeneracy at X and Y, marked by the red circle and (b) with $J_{1}=-0.04J$, a gap opens up at X and Y.
@@
$\hskip-5.69046pt\hbar\omega^{\pm}_{\bm{q}}=\frac{1}{\sqrt{2}}\sqrt{\left(A^{11}_{\bm{q}}\right)^{2}+\left(A^{22}_{\bm{q}}\right)^{2}-2\left(B^{12}_{\bm{q}}\right)^{2}+2\left|A^{12}_{\bm{q}}\right|^{2}\pm\mathcal{E}^{2}_{\bm{q}}}\,$, (6)
@@
[…]
@@ V Topology of the magnon bands: Berry curvature and Chern numbers
Figure 3: Berry connections ${\cal\bm{A}}^{\pm}_{\bm{q}}$ [Eq. (9)] of the upper and lower branches for the AM coupling $J_{1}=-0.04J$, DMI constant $D=0.5J$, and Zeeman field $H=H_{s}/8$.
@@
and show it across the first Brillouin zone for the two branches in Fig. 4 for different values of the AM coupling. The $C_{4}$ lattice rotation combined with a $C_{2}$ spin-rotation symmetry about the $\bm{\hat{z}}$ axis of the Hamiltonian [Eq. (1)] is reflected in the Berry curvature as $\Omega^{\pm}_{C_{4}\bm{q}}=\Omega^{\pm}_{\bm{q}}$, as evident in Fig. 4. Applying this symmetry twice successively results in the even-under-inversion property of the Berry curvature, i.e., $\Omega^{\pm}_{-\bm{q}}=\Omega^{\pm}_{\bm{q}}$. Finally, the topology of the magnon bands in two-dimensional systems is characterized by the (first) Chern number defined in terms of the Berry curvature as
$C^{\pm}=\frac{1}{2\pi}\int_{\rm FBZ}\Omega_{\bm{q}}^{\pm}\,\,dq_{x}\,dq_{y},$ (12)
where the integration is performed over the first Brillouin zone (FBZ). Evaluating this expression for the two gapped bands yields $|C^{\pm}|=1$  [^1] The integral in Eq. (12) is performed numerically using the analytical form of the Berry curvature for a wide range of parameter sets. with $C^{+}=-C^{-}$, thereby clearly demonstrating that the magnon bands in our model are topologically non-trivial.
@@
Figure 4: Evolution of the Berry curvature [Eq. (11)] of the upper (top row) and lower (bottom row) magnon branches as the absolute value of the AM coupling $J_{1}$ decreases from left to right. The other parameters are fixed at the DMI constant $D=0.5J$ and the Zeeman field $H=H_{s}/8$.
We find the following interesting properties of the Berry curvature and Chern number for each magnon branch. Under the sign reversal of either $H$ or $D$, the spin-wave matrix $\bm{M}_{\bm{q}}$ (see Appendix A for details) is complex conjugated, $\bm{\bar{M}}_{\bm{q}}$. Under complex conjugation, the eigenvalue equation for the magnon branches, $\bm{\eta}\bm{M}_{\bm{q}}|\Psi^{\pm}_{\bm{q}}\rangle=E^{\pm}_{\bm{q}}|\Psi^{\pm}_{\bm{q}}\rangle$ [Eq. (76) in Appendix A] becomes $\bm{\eta}\bm{\bar{M}}_{\bm{q}}|\bar{\Psi}^{\pm}_{\bm{q}}\rangle=E^{\pm}_{\bm{q}}|\bar{\Psi}^{\pm}_{\bm{q}}\rangle$. This shows that flipping the sign of $H$ or $D$ leaves the magnon energy eigenvalues unchanged, while the eigenvectors are complex conjugated, $|\bar{\Psi}^{\pm}_{\bm{q}}\rangle$. Consequently, the Berry connection changes sign since $\langle\Psi_{\bm{q}}^{\pm}|\bm{\eta}\,\bm{\nabla}_{\bm{q}}|\Psi_{\bm{q}}^{\pm}\rangle$ is purely imaginary. Hence, the Berry curvature [Eq. (11)] changes as $\Omega^{\pm}_{\bm{q}}\to-\Omega^{\pm}_{\bm{q}}$, resulting in $C^{\pm}\to-C^{\pm}$. When the AM coupling is reversed, $J_{1}\to-J_{1}$, the Berry curvature also changes sign at each wave vector, but unlike the previous cases, its absolute value does not remain unchanged. In this case, there is no simple relation between the old and new eigenvectors. Nevertheless, the change happens in such a way that we still find $C^{\pm}\to-C^{\pm}$.
@@ VI Thermal Hall conductivity
As a physical consequence of the nontrivial topology of the magnon bands, we consider the magnon thermal Hall effect, i.e., the effect of heat transfer by magnons in the direction perpendicular to the applied temperature gradient [19, 20, 21]. This phenomenon has been studied extensively in both ferromagnetic [19, 21, 65, 22, 23, 24, 66, 31, 32, 33, 35, 39, 40, 41, 42] and antiferromagnetic [62, 43, 67, 44, 45, 46, 47, 48, 50, 51, 68, 49] systems, while its realization in an altermagnetic system has only recently begun to attract attention [69]. To compute the thermal Hall conductivity, we use the formula [19, 20, 65]
@@
The Zeeman field dependence of $\kappa_{xy}$ is shown in Fig. 6. As the field is reversed, $\kappa_{xy}$ changes sign and, notably, exhibits a sharp jump across zero at low temperatures. This non-monotonic behavior of $\kappa_{xy}$ for low fields at low temperatures follows from the field dependence of the Berry curvature $\Omega^{-}_{\bm{q}}$ of the lower magnon branch near $\Gamma$ point, see the inset in Fig. 6. Note that the magnons in the lower branch from the vicinity of $\Gamma$ point mainly contribute to $\kappa_{xy}$ at low temperatures.